### v3

# Which Transformer to Favor: A Comparative Analysis of Efficiency in Vision Transformers

Abstract
Self-attention in Transformers comes with a high computational cost because of their quadratic computational complexity, but their effectiveness in addressing problems in language and vision has sparked extensive research aimed at enhancing their efficiency. However, diverse experimental conditions, spanning multiple input domains, prevent a fair comparison based solely on reported results, posing challenges for model selection. To address this gap in comparability, we perform a large-scale benchmark of more than 45 models for image classification, evaluating key efficiency aspects, including accuracy, speed, and memory usage. Our benchmark provides a standardized baseline for efficiency-oriented transformers. We analyze the results based on the Pareto front – the boundary of optimal models. Surprisingly, despite claims of other models being more efficient, ViT remains Pareto optimal across multiple metrics. We observe that hybrid attention-CNN models exhibit remarkable inference memory- and parameter-efficiency. Moreover, our benchmark shows that using a larger model in general is more efficient than using higher resolution images. Thanks to our holistic evaluation, we provide a centralized resource for practitioners and researchers, facilitating informed decisions when selecting or developing efficient transformers.  [^1] https://github.com/tobna/WhatTransformerToFavor
1 Introduction
Figure 1: Pareto front (dotted line) of throughput and accuracy. Marker shapes and hues signify the different overarching strategies for making ViT more efficient. Pareto optimal models have a black dot, while the others have a white dot. The size of markers signifies if we fine-tuned using the standard $224$px or the larger $384$px resolution. We mark the baselines ViT-Ti@224 (A), ViT-Ti@384 (B), and ViT-S@224 (C) and also plot the Pareto front for high resolution images (dashed line).
The introduction of the Transformer [63] prompted its widespread adoption in both language and vision communities. In particular for image classification, the Vision Transformer (ViT) [16] has positioned itself as one of the best known applications of the original architecture, outperforming traditional CNN architectures on benchmarks like ImageNet [76, 75]. However, a major challenge in working with Transformer models is dealing with the computational complexity of the self-attention mechanism. This mechanism enables the Transformer to capture global dependencies between pairs of tokens, but it has a computational complexity of $\mathcal{O}(N^{2})$ in the input length $N$, making it impractical for long sequences and high-resolution images.
Efforts have been made to reduce the computational load of Transformer models, particularly for resource-constrained settings, such as embedded systems [64, 51]. Researchers have explored numerous strategies, like implementing sparse, local, or kernelized attention mechanisms, as well as token removal criteria to decrease the sequence length. However, selecting the most efficient model that meets certain performance standards remains a challenging task. This task becomes particularly difficult, knowing that “efficiency” can refer to different concepts, like training resources, inference speed, or memory requirements. Moreover, difficulties identifying the most efficient, yet highly performant vision transformers are accentuated by different training and evaluation conditions reported in the literature. It is even more unclear how tradeoffs in efficiency for models proposed for NLP carry over to vision tasks.
To address these problems of fairly comparing efficient ViTs, we design an impartial test bed, training models from scratch on the same data, using the same setup. While many papers at most report inference speed [41, 37, 38, 61], our benchmark enables us to fairly provide empirical values across multiple dimensions of efficiency. We identify common strategies employed to enhance model efficiency and utilize them to correlate changes in efficiency metrics with the respective strategies. Through our benchmark, we provide a centralized resource of comprehensive baselines in the domain of image classification, conducting a thorough review of the current state of research on efficient Transformers. In particular, we provide baselines for Transformers proposed for NLP, which can be trivially adapted to CV.
We analyze model speed and memory requirements both at training and inference time, as well as parameter efficiency. This reveals that not all efficiency claims are realized. Our analysis is based on the Pareto front, the set of models that provide an optimal tradeoff between model performance and one aspect of efficiency. It lets us analyze the complex multidimensional tradeoffs involved in judging efficiency. We plot inference throughput against accuracy in Figure 1 to visualize the Pareto front. Here, a model is Pareto optimal (black dot) if and only if there is no model that is both more accurate and faster at the same time. Through this analysis, we show that ViT remains Pareto optimal across three of the four empirical metrics we track, and also find other efficiency strategies that are Pareto optimal for different metrics. This benchmark, based on more than 200 experiments, provides researchers and practitioners with a valuable resource, aiding them in choosing the most efficient and effective model architecture for their specific use case.
Main Results
  • A fair large-scale benchmark for efficiency in transformer models run on 45+ models, closing the comparability gap due to different experimental conditions.
  • We find that a well-trained ViT still is Pareto optimal for three out of four metrics, challenging claims of other models being more efficient [68, 43, 10].
  • Our analysis uncovers novel findings like that in most cases, scaling up the model size is more efficient than scaling up the image resolution.
2 Related Work
This section provides an overview of relevant work on transformer surveys and approaches to measure efficiency, while Section 3.3 delves into specific efficient architectures.
Surveys on Efficient Transformers
Efficiency has become a critical aspect of transformers, leading to the creation of multiple surveys on efficient transformers providing valuable taxonomies and insights. [55] focuses on efficiency gains primarily in NLP, while [19] and [81] collect general approaches for enhancing model efficiency, including those applicable to transformers. [41, 35] present an extensive list of efficient ViTs, classified based on aspects such as computational complexity, robustness, and transparency, comparing the efficiency in terms of parameters, ImageNet accuracy, and inference speed, using information from the original papers. Surveys on ViT-like models [22, 72, 26] focus on categorizing models according to different vision tasks, while [82] specializes in dense prediction. Specialized surveys investigate the application of transformers in specific domains, like action recognition [60], image restoration [2], medical imaging [50, 31, 23, 40], remote sensing [1], speech recognition [29], language processing [9], time series [66], or multimodal tasks [69].
We build on this by analyzing efficiency gains in terms of broader strategies for efficiency. In contrast to previous work, we close the gap of different training strategies and enable a fair comparison of efficient architectures under the same conditions based on more than new 200 experiments.
Evaluation of Efficiency
Efficiency evaluation in deep learning models is another area of investigation. [4, 14] provide an overview of efficiency aspects and metrics, along with measurement methodologies, highlighting the potential pitfalls of relying solely on theoretical metrics. [8] conducts a highly regarded survey on the efficiency of CNNs, and [34] compares the efficiency of their novel architectures with older models using the Pareto front of throughput and accuracy. Finally, [54] introduces a benchmark on synthetic data to quantify transformer model performance on long sequences in NLP, evaluating 10 models.
We extend previous work by evaluating more than 45 efficient Transformers for CV, measuring empirical metrics on the same hardware and using real data. The in-depth analysis of our benchmark’s results reveals new insights.
3 Efficient Transformers for Vision
First, we present our methodology for efficiency assessment. Then we briefly describe ViT, as well as a diverse collection of efficient transformers that have been proposed and what these change compared to ViT. These models from the domains of CV and NLP all claim to improve the efficiency of the baseline transformer. We systematically select efficient transformers based on popularity, diversity, and novelty of their approaches. See the supplementary material for more detailes on model selection.
3.1 Quantifying Efficiency
The term efficiency can take on different meanings; it is therefore crucial to consider multiple dimensions when evaluating a model’s efficiency. In this paper, we analyze model speed and memory requirements at training and inference time and contrast them with accuracy.
Due to the complex multidimensional tradeoffs, efficiency cannot be captured by a single number. Hence, we focus on the Pareto front – the boundary of the landscape of efficient models – to identify models that achieve the best trade-offs between two metrics. The Pareto front represents the set of models that take the best compromises between two metrics. A Pareto optimal model outperforms every other model in at least one of the two metrics. See the Pareto front in Figure 1, where for each model that is not on the Pareto front (white dot) there is a model on the Pareto front (black dot) that is both faster and more accurate.
3.2 Core Elements of ViT
[…]
$A=\operatorname{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{\text{head}}}}\right)\in\mathbb{R}^{N\times N},$
where the softmax is calculated row-wise. This matrix encodes the global interactions between every possible pair of tokens. It is also the reason why the attention mechanism has an inherent computational (space & time) complexity of $\mathcal{O}(N^{2})$. The output sequence of the attention-mechanism is a weighted sum of the input sequence, using the value matrix ($V$):
$\begin{split}X_{\text{out}}=AV=\operatorname{softmax}\left(\frac{QK^{\top}}{\sqrt{d_{\text{head}}}}\right)V.\end{split}$ (1)
[…]
Figure 2: List of efficient Transformers (citation key in brackets) categorized at two levels: 1. Where does the approach change ViT? 2. How does the approach change ViT?
3.3 Efficiency-Improving Changes
After having gone over the backbone that constitutes a ViT, we discuss the most important modifications, which have been proposed to make it more efficient. In order to analyze the overarching strategies that are taken, we classify the efficient models systematically using a two-step approach. First, we identify the specific components — the token mixing mechanism, the token sequence, or the MLP block — of the baseline ViT architecture that are modified (i.e., “where”), as visualized in Figure 2. On the second level, we classify the different strategies to enhance the efficiency (i.e., “how”). While this taxonomy is not meant to be a comprehensive overview of ViT-like models, it is proposed as a tool for identifying the best and most popular strategies to make vision transformers more efficient. Note that our categorization scheme aligns closely with taxonomies that have been recently proposed [19, 55, 47].
(i) Token Mixing
The first and most popular approach is to change the token mixing mechanism, which directly tackles the $\mathcal{O}(N^{2})$ computational complexity of self-attention. There are multiple strategies through which this can be accomplished. Some methods approximate the attention mechanism with reduced computation, which can be achieved by matrix decomposition, changing the order of operations, or fixing attention values. Other approaches combine attention with CNNs to perform sub-sampling in the attention mechanism or reduce the number of uses of the attention mechanism. Finally, some methods discard the attention mechanism and instead introduce entirely new token mixing strategies.
Low-rank attention leverages the fact, that $Q$, and $K$ in Equation 1 are matrices of shape $N\times d$, which makes ${QK^{\top}\in\mathbb{R}^{N\times N}}$ a matrix of rank $r\leq d\ll N$. The Linformer [65] utilizes this to project the sequence direction of $K$ and $V$ down to dimension $k\ll N$, without reducing the informational content of the attention matrix too much. Similarly, the Nyströmformer [68] uses the Nyström method of matrix decomposition to approximate the matrix $QK^{\top}$. The approximate attention mechanism’s output is then computed with linear complexity by applying the $\operatorname{softmax}$ to each part individually. The approach of XCiT [17] utilizes a transposed attention mechanism:
$Y^{\top}=\operatorname{softmax}\left(\frac{1}{\sqrt{d}}Q^{\top}K\right)V^{\top}.$
Here, $Q^{\top}K\in\mathbb{R}^{d\times d}$ is used to replace the low-rank matrix $QK^{\top}$ to define the globally informed filter ${\operatorname{softmax}\left(\frac{1}{\sqrt{d}}Q^{\top}K\right)\in\mathbb{R}^{d\times d}}$ which is applied to each token individually. Since both $Q$ and $K$ are likely to be of rank $d$, the former most likely has full rank, which enables more efficient information encoding.
Sparse attention is an alternative approach to these dynamic approximations of global interactions. Normal attention tends to focus on only a few input tokens, giving low scores to most others [27]. Sparse attention takes advantage of this by fixing most of the attention weights at zero and only calculating the most important ones. The parts of the attention matrix, which are determined dynamically, can be based on a fixed pattern, like in the widely used Swin Transformer [37], and SwinV2 [36], where attention is performed only inside local sets of image tokens, or in HaloNet [62], where each token can only attend to its neighbors. Alternatively, Routing Transformer [45] determines sets of tokens that exchange information by grouping them based on their content. In contrast, Sinkhorn Transformer [53] fixes local groups and lets tokens only attend to a different group using a dynamic permutation matrix. Informer [80] only calculates attention values for the $c\cdot\ln N$ most important tokens. Another way of keeping the attention matrix sparse, is by sub-sampling the keys and values. For example, Wave-ViT [73] uses a discrete wavelet transform for sub-sampling.
Fixed attention is an extreme example of setting attention values beforehand, which then only depend on the tokens positions. This is explored in the Synthesizer [52].
Kernel attention is an approach that changes the order of computations in Equation 1. Instead of using the softmax on the product $QK^{\top}$, a kernel $\varphi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is applied to the queries $Q$ and keys $K$ individually:
$Y=\varphi(Q)\varphi(K)^{\top}V,$
which can be calculated with linear complexity $\mathcal{O}(N)$ in the number of tokens. Various kernels have been proposed, including the random Gaussian kernel, introduced in Performer [13] and expanded on in the FourierLearner-Transformer [12] by adding a relative positional encoding in phase space. Scatterbrain [11] combines this attention mechanism with sparse attention. Poly-SA [3] uses the identity function $\varphi\equiv\mathrm{id}$, Linear Transformer [25] uses an ELU, and SLAB [21] uses a ReLU kernel. Hydra ViT [6] uses an L2-normalization kernel and employs $H=d$ heads.
Hybrid attention combines convolutions with the attention mechanism. EfficientFormerV2 [33] uses convolutions initially to focus on local interactions, and then employs the attention mechanism to capture global interactions, while EfficientViT [7] utilizes convolutions first and kernel attention, using the ReLU kernel later on. Next-ViT [32] alternates $N_{\text{Conv}}\in\mathbb{N}$ convolution blocks with a single attention block. Another approach is to use convolutions inside the attention mechanism to create locally informed queries, keys, and values, as in CvT [67] and ResT [77], where it is used to subsample keys and values, and in CoaT [70].
Fourier attention uses the Fast Fourier Transform (FFT) in the calculation of the attention mechanism to reduce the complexity to $\mathcal{O}(N\log N)$. FNet [30, 48] directly exploits the FFT for $\mathcal{O}(N\log N)$ complexity token mixing, while GFNet [43] utilizes it for a global convolution and AFNO [20] uses an MLP in Fourier space.
Non-attention shuffling refers to techniques of capturing token interactions without using attention. For example MLP-Mixer [56] uses a fully connected layer for global interactions, while FastViT [61] mixes tokens using depth-wise convolutions. EfficientMod [38] modulates a convolutional context with a value matrix and FocalNet [71] modulates a hierarchy of contexts extracted by convolutions. SwiftFormer [49] calculates a global query to multiply with the keys and also replaces the MLP block by a CNN.
(ii) Token Sequence
The second category, token sequence, is more prevalent in efficient transformers used in CV compared to NLP. The idea is to remove redundant information typically contained in images, and in doing so, reducing computational costs without significantly affecting the model’s performance. Methods aim to reduce the token sequence by removing unnecessary image patches, such as the background, by merging tokens to minimize redundant information, or by summarizing the information into a smaller number of abstract tokens representing higher-level information. These models leverage the $\mathcal{O}(N^{2})$ complexity of the self-attention mechanism to attain a large reduction in computational cost, as removing $30\%$ of the tokens reduces the operations needed by around $50\%$. While some of these methods can directly be applied to trained models, performance improves when implementing them during training [5], which we do.
Token Removal requires identifying which tokens to remove without losing critical information. To that end, Dynamic ViT [42] uses the Gumbel-softmax for token retention probabilities, and A-ViT [74] learns halting probabilities that weight the token outputs of different depths. In contrast, EViT [34] avoids introducing extra parameters by utilizing the previous layer’s attention matrix.
Token merging is another strategy to remove redundant information. While a version of EViT [34] merges the unimportant tokens, ToMe [5] merges tokens based on their similarity, using a fast bipartite matching algorithm.
Summary tokens, condensing the token sequence into a few new tokens, contrasts the approaches so far, which modify the existing tokens. CaiT [59] uses token summary via cross-attention on a single token to gather global information on the classification decision in the last layers. Token Learner [46] creates a set number of summary tokens by employing dynamic sums over the image tokens, and STViT [10] initializes summary tokens with local information by using strided convolutions, then injecting global information into those using cross-attention.
(iii) MLP Block
The final way in which proposed methods change the architecture of transformers is by moving compute to the MLP block, which has linear complexity with respect to the sequence length. Despite room for efficiency gains, we have identified only two methods taking this approach, which we include to make this benchmark more diverse.
More MLPs can be used to tackle this problem by introducing multiple parameter sets for each MLP block and passing different tokens through blocks with different parameters. Switch Transformer [18] can have multiple times the number of parameters of the ViT without introducing lots of extra computations. In contrast, HiViT [78] exchanges early attention layers for additional MLP blocks.
Out categorization offers a structured framework for understanding approaches of improving ViT efficiency. This enables us to compare broad trends across strategies.
4 Experimental Design
We conduct an extensive series of over 200 experiments on the more than 45 models described in Section 3.3 to evaluate which models are ideal under a given set of constraints. We measure efficiency w.r.t. the full models rather than individual attention modules, since methods that reduce the token sequence (Section 3.3 (ii)), and others, like XCiT, require additional modifications to the architecture, which we include. Note, however, that many models only introduce changes in the attention mechanism, effectively turning our benchmark into a module-level comparison for these.
4.1 Baselines
To find the most efficient models, we need to quantify their efficiency gains by comparing them to the original ViT [16] and its follow-up versions [57, 58]. We also include ResNet50 [24] in our evaluation as a representative baseline for CNN architectures and a point of reference across papers. All evaluations are based on the ImageNet-1k dataset for classification [15] as it is one of the most commonly used datasets in this domain. Results on other datasets (Stanford Cars [28], Oxford Flowers 102 [39], MIT Places365 [79]) can be found in the supplementary material.
4.2 Training Pipeline
We compare models on even grounds by training with a standardized pipeline. To reduce bias, our pipeline is relatively simple and only consists of elements commonly used in CV. In particular, we refrain from using knowledge distillation to prevent introducing bias from the choice of teacher model. Any orthogonal techniques, like quantization, sample selection, and others, are not included as they can be applied to every model and would manifest as a systematic offset in the results. To avoid issues from limited training data, we pre-train all models on ImageNet-21k [44].
Models are trained for a total of 140 epochs using the pipeline introduced in [58]. This is an updated version of DeiT [57], which has been adopted throughout the literature for training efficient ViTs (see Table 1). As this pipeline only employs standard CV practices, we consider it a fair point of comparison for evaluating all models.
Pretraining is conducted on ImageNet-21k for 90 epochs at resolutions of $224$ and $192$ pixels, followed by fine-tuning on ImageNet-1k at $224$ and $384$ pixels for 50 epochs. In cases where model training is unstable, we default back to the hyperparameters reported in the corresponding publications. All training is conducted using 4 or 8 NVIDIA A100 GPUs. While most of the models work well with our training pipeline, some did not converge. See the supplementary material for more details.
4.3 Efficiency Metrics
As we saw, efficiency has multiple dimensions, and therefore we need to capture a variety of metrics. For comparison with other papers and to evaluate their use as proxy metrics for speed and memory requirements, we capture the theoretical metrics of number of parameters and FLOPS. These are used by the community to provide estimates of the model’s representational capacity and computational requirements [41, 55, 43]. However, theoretical metrics do not always correlate with real-world performance, especially when comparing different architectures [14, 4]. Instead, empirical metrics, obtained by running models on hardware, offer more informative evaluations, which is why we track GPU time for training, inference throughput at optimal batch sizes, and VRAM requirements. While these empirical metrics can be sensitive to hardware and software configurations, we ensure consistent evaluations by employing the same setup for all models. We evaluate the variance of our metrics in the supplementary material.
5 Results
Table 1: ImageNet-1k accuracy of the original papers and the new training pipeline. ✓ marks original pipelines based on DeiT [57]. Models are trained on $224$px images, unless marked with $\uparrow$ ($384$px) or $\downarrow$ ($112$px). Results with D use knowledge distillation and E uses exponential moving average.
Model Original Ours
DeiT Accuracy Accuracy
ViT-S (DeiT) 79.8 82.54
ViT-S (DeiT III) 82.6 82.54
XCiT-S 82.0 83.65
Swin-S 83.0 84.87
SwinV2-Ti 81.7 83.09
Wave-ViT-S 82.7 83.61
Poly-SA-ViT-S 71.48 78.34
SLAB-S 80.0 78.70
EfficientFormer-V2-S0 75.7D 71.53
CvT-13 83.3$\uparrow$ 82.35
CoaT-Ti 78.37 78.42
EfficientViT-B2 82.7$\uparrow$ 81.52
NextViT-S 82.5 83.92
ResT-S 79.6 79.92
FocalNet-S 83.4 84.91
SwiftFormer-S 78.5D 76.41
FastViT-S12 79.8$\uparrow$ 78.77
EfficientMod-S 81.0 80.21
GFNet-S 80.0 81.33
EViT 79.4 82.29
DynamicViT-S 83.0D 81.09
EViT Fuse 79.5 81.96
ToMe-ViT-S 79.42 82.11
TokenLearner-ViT-8 77.87$\downarrow$ 80.66
STViT-Swin-Ti 80.8 82.22
CaiT-S24 82.7 84.91
5.1 Improved Training Pipeline
To validate the fairness of our training pipeline, in Table 1 we compare ImageNet-1k accuracy reported in the original papers (whenever reported) with the one we obtained to see if there are any obvious outliers. Note, that a substantial fraction (13 out of 26) of papers explicitely base their training pipelines on DeiT [57], making them a good fit for training with the updated version in DeiT III [58]. We observe that almost all models trained with this pipeline achieve a higher accuracy; 0.85% on average, with the largest gain being +6.86% for Poly-SA. The models reporting higher performance than what we obtain use knowledge distillation, which we avoid to not induce bias, or a higher resolution, which we find is inefficient.
Most categories contain models that achieve peak accuracy of approximately 85% (see the supplementary material). However, two notable exceptions are observed: the Kernel Attention class, which might pose challenges in optimizing models, and the Fixed Attention class, where the use of a constant attention matrix inherently results in a lower accuracy. Overall, our selection of the training pipeline as a reference for a fair comparison is supported by the improvements in results compared to the original papers, regardless of architectural differences.
Figure 3: Accuracy (red line, right y-axis) and accuracy per parameter (bars, left y-axis) of models ordered by accuracy at a resolution of $224$px. 14 of 47 models of intermediate size are grouped into others. See the supplementary material for the full plot.
5.2 Number of Parameters
The number of parameters is used in the literature as a proxy metric for tracking the model complexity and the overall computational cost of using a model. When analyzing the parameter efficiency of different models (Figure 3), it is evident that for most smaller models, the accuracy per parameter remains relatively constant at a baseline level of about $4\times 10^{-8}\frac{\%}{\text{param.}}$. This value is approximately halved for larger models, indicating diminishing returns on scaling the model size. For ViT, there is a noticeable drop in accuracy per parameter across the model sizes. While ViT-Ti outperforms models of comparable accuracy, and ViT-S performs on par with similar models, ViT-B slightly underperforms when compared to other, larger models. Significant outliers are the smaller Hybrid Attention models EfficientFormerV2-S0 and CoaT-Ti, which exhibit the highest accuracy per parameter, as well as those Non-attention Shuffling models which incorporte convolutions, like SwiftFormer and FastViT. In the MLP category, we find HiViT to have significantly higher and Switch ViT with significantly lower accuracy per parameter than surrounding models. This suggests that the combination of attention and convolutions allows for the development of highly parameter-efficient models.
5.3 Speed
Figure 4: Pareto front of finetuning time and accuracy for models which need less than 50 hours for finetuning. We include the full plot in the supplementary material.
Whether driven by strict requirements for real-time processing or the desire to obtain model outputs within reasonable timeframes, inference speed directly impacts the usability of deployed models. The models we evaluate often claim a superior throughput vs. accuracy trade-off compared to ViT. However, our evaluation in Figure 1 reveals that ViT remains Pareto optimal at all model sizes. Only few models, namely Synthesizer-FR, NextViT, and some Sequence Reduction models, show improvements in the Pareto front when compared to a ViT of comparable size.
We also analyze the Pareto optimality with models finetuned on Stanford Cars [28], Oxford Flowers 102 [39], and MIT Places365 [79] in the supplementary material, observing a high Spearman correlation between the accuracy of different datasets. In particular, the Pareto optimal models of throughput and accuracy are largely the same, which is a strong indicator for the generalizability of our results. Even when assessing throughput on a CPU instead of a GPU, we observe only a high Spearman correlation with only few outliers. Again, we find the Pareto front to be robust to the choice of hardware. See the supplementary material for more information about these results.
In the Pareto front of fine-tuning time in Figure 4, we see many similarities to the one of inference time. Here, some of the token sequence based models are highly efficient. In particular, TokenLearner emerges as a standout performer with the fastest fine-tuning speed while achieving a competitive accuracy of 77.35%. For larger models, we find ViT-B and FocalNet (-Ti and -S) to be Pareto optimal.
Generally, ViT is still a solid choice for speed, with some Token Sequence models trading off accuracy for speed.
Figure 5: Pareto front (dotted line) of training memory at our default batch size of $2048$ (left) and inference memory at the minimum batch size of $1$ (right) and accuracy for models with less than 225GB of VRAM for training and 1.25GB for inference.
5.4 Memory
Figure 5 (right) reveals that the tradeoffs when optimizing for VRAM usage during inference are very different from the ones for speed or training VRAM (left side), making it the only metric where ViT is not Pareto optimal. At inference time, the Hybrid Attention models CoaT, CvT, ResT, and NextViT stand out from the rest. Additionally, CaiT, and EViT@384px for the larger sizes and EfficientMod for the smaller ones are Pareto optimal. This observation suggests that, similar to the number of parameters, Hybrid Attention models excel in low memory environments.
In contrast, the aspect of training memory (Figure 5; left) exhibits a similar pattern as observed in throughput. Here, CoaT, CvT, and ResT need more memory than other models of comparable accuracy and again ViT is Pareto optimal, along with some Sequence Reduction models and Synthesizer-FR. NextViT is the only model that is Pareto optimal for both training and inference memory. When training at a resolution of $384$px most models need more than $225$GB of VRAM and therefore are not plotted.
Overall, inference memory is the most different from the other Pareto fronts with Hybrid Attention models excelling.
5.5 Scaling Behaviors
Our observations reveal that fine-tuning at a higher resolution is inefficient. While it may result in improved accuracy, it entails a significant increase in computational cost, leading to a substantial reduction in throughput. In turn, scaling up the model ends up being more efficient. This is shown in Figure 1, where the best models using high resolution images are two to four times slower than a low resolution model at the same accuracy (compare Pareto front and HR Pareto front). Consequently, no model fine-tuned at the higher resolution is Pareto optimal. For example, when scaling from tiny to small (A to C in Figure 1), we gain a lot more accuracy for the lost speed than when scaling up the image size (A to B). For training memory, this trend is even more pronounced (see HR Pareto front in Figure 5), with most of the high resolution models requiring more than $225$GB of VRAM for training and thus being off the chart. Overall, many papers focus on reducing the $\mathcal{O}(N^{2})$ complexity of self-attention to deal with long sequences, however sequences in image classification are typically not long enough to show the theoretical gains. On the contrary, it is more efficient to use shorter sequences instead. We believe that efficient mechanisms focusing on these shorter sequences are a fruitful area for future research.
5.6 Correlation of Metrics
Table 2: Correlation between the number of floating point operations, number of parameters ($\theta$), fine-tuning time, training memory, inference time, and inference memory.
$\operatorname{corr}(x,y)$ $\theta$ Training Inference
Time Mem Time Mem
FLOPS 0.30 0.72 0.85 0.48 0.42
$\theta$ 0.05 0.18 0.02 0.40
Training Time 0.89 0.81 0.17
Training Mem 0.71 0.48
Inference Time 0.13
We run a correlation analysis between the efficiency metrics in Table 2. The highest correlation of $0.89$ is between fine-tuning time and training memory. This suggests a common underlying factor or bottleneck, possibly related to the necessity of memory reads during training. Understanding this relationship can provide valuable insights into the factors influencing training efficiency. Intriguingly, the highest correlation coefficient involving a theoretical metric is $0.85$ between FLOPS and training memory, suggesting the possibility for a rough estimation. For other metrics, we reaffirm the limited reliability of estimating computational costs solely based on theoretical metrics [14, 4]. Consequently, assessing model efficiency in practice requires the empirical measurement of throughput and memory requirements.
6 Discussion & Conclusion
WhichTransformerto Favor?1. MainConstraint2. Phase3. Transformers to FavorSpeedMemoryInferenceTrainingInferenceTraining{ViT NextViT, ToMe, TokenLearner, }{ViT, TokenLearner, FocalNet}{ViT, NextViT, CoaT, CvT, }{ViT, ToMe, TokenLearner, NextViT, …}
Figure 6: Suggestions on Which Transformer to Favor in image classification. First, choose the main constraint (1.), speed, or memory, and the phase you want to optimize this constraint at (2.): Inference or training. Based on that, we provide recommendations on which transformer to favor (3.) from the taxonomy classes of Baseline, Sequence Reduction, Hybrid Attention, Non-Attention Shuffling and Sparse Attention. See the supplementary material for a full list of Pareto optimal models for each metric.
In this paper, we provide a thorough benchmark of efficiency oriented transformers for vision, facilitating a fair, competitive comparison of contributions from NLP and CV and revealing some unexpected trends. Our analysis shows ViT’s continued Pareto optimality across multiple metrics, reaffirming its effectiveness as a baseline, despite other models claiming to be more efficient. Results also prove that scaling the model size is, in most cases, more efficient than using higher resolution images. This goes against the trend of efficient models being evaluated using high resolution images [59, 37, 17, 61]. Comparing different efficiency metrics, we reaffirm the limitations of estimating computational costs solely based on theoretical metrics.
When addressing the question of which transformer to favor, our benchmark offers actionable insights in the form of models and strategies to use (see Figure 6). ViT remains the preferred choice overall. However, Token Sequence methods can become viable alternatives when speed and training efficiency are of importance. For scenarios with significant inference memory constraints, considering Hybrid CNN-attention models can prove advantageous.
Looking ahead, we posit that the establishment of this comparable, fair benchmark and analysis will drive substantial advancements in efficiency-oriented transformers for CV, benefitting researchers and practitioners making architectural decisions for resource-constrained environments.
Acknowledgements
This work was funded by the Carl-Zeiss Foundation under the Sustainable Embedded AI project (P2021-02-009).